OpenStudy (anonymous):

Decide whether each equation has one solution, no solutions, or infinitely many solutions. 2(x – 3) = 2x one solution no solutions infinitely many solutions 3(y – 3) = 2y – 9 + y one solution no solutions infinitely many solutions

OpenStudy (anonymous):

wait... not reverswe haha i meant just regular.

OpenStudy (anonymous):

Haha ok! :)

OpenStudy (anonymous):

whatcha get?

OpenStudy (anonymous):

2x - 6 = 2x

OpenStudy (skullpatrol):

Both have no solutions.

OpenStudy (anonymous):

okay combine like terms.

OpenStudy (anonymous):

and yu get?

OpenStudy (anonymous):

2x-6

OpenStudy (anonymous):

OR WAIT 4x - 6?

OpenStudy (anonymous):

no yu suppose to subtract 2x from both sides.

OpenStudy (anonymous):

Im stuck

OpenStudy (anonymous):

I keep getting -6

OpenStudy (anonymous):

okay 2x-6=2x -2x -2x ---------- -6=0

OpenStudy (anonymous):

ok...so it has no solution.

OpenStudy (anonymous):

right. and then yu do the smae to the next one.

OpenStudy (anonymous):

@skullpatrol what about these two: 10x – 2 – 6x = 3x – 2 + x one solution no solutions infinitely many solutions 4(x + 3) + 2x = x – 8 one solution no solutions infinitely many solutions

OpenStudy (anonymous):

@hereweg0 my mom did the next one lol.

OpenStudy (anonymous):

okay nice.

OpenStudy (anonymous):

Ok I am gonna open up a new Q for the next 2.

OpenStudy (anonymous):

Closing this one...